# Can GPT models Follow Human Summarization Guidelines? A Study for Targeted Communication Goals

### v2

@@ Appendix E Output Length
In Figure 4 and Figure 5, we present statistics on the number of words in summaries generated using different prompts on the DialogSum and DECODA datasets compared to reference summaries, respectively.
The summaries generated by ChatGPT are all longer than the reference summaries. The use of WordLimit prompt showed a shorter average output length, and its use as a second step makes the summaries generated with HG(R) more synthetic (HGRWL).
As for GPT-4, it better respects the instructions regarding output length, we can see that the distribution of 4-WL and 4-HGRWL is more restricted in both figures, while 4-HGR exceeds the length of the reference summaries.
@@ F.1 DialogSum Discrepancy Data Points
Table 15 presents the DialogSum cases where GPT-generated summaries exhibit the most significant scoring discrepancies (low ROUGE-L but high BERTScore values) for 4-HGRWL.
@@ F.2 DECODA Discrepancy Data Points
Table 16 displays English translations of DECODA cases where GPT-generated summaries exhibit the most significant scoring discrepancies (low ROUGE-L but high BERTScore values) for 4-HGRWL.
@@
Tables 17 and 18 present the original French DECODA cases where GPT-generated summaries exhibit the most significant scoring discrepancies (low ROUGE-L but high BERTScore values) for 4-WL and 4-HGRWL respectively, as analyzed in Section 3.3.
@@ Appendix G Datasets: License or Terms of Use
The DialogSum corpus we used in this study is composed of resources freely available online without copyright constraint for academic use. The training and validation sets of the DECODA dataset can be downloaded from their website131313https://pageperso.lis-lab.fr/benoit.favre/cccs/, upon acceptance of the corresponding usage and sharing terms, and the test set is available only on request from the authors.

### v3

@@ Appendix E Output Length
Figure 7 and Figure 8 present summary length statistics for different prompts on the DialogSum and DECODA datasets, respectively, compared to reference summaries.
The summaries generated by ChatGPT are consistently longer than the reference summaries. Applying the WordLimit prompt reduced the average output length, and using it as a second step after the HG(R) prompt produced more concise summaries (HGRWL).
Regarding GPT-4, it demonstrated better adherence to length instructions. The distributions for 4-WL and 4-HGRWL are more constrained in both figures, while 4-HGR exceeds the reference summary length.
@@ F.1 DialogSum Discrepancy Data Points
Table 15 presents DialogSum cases with the most significant scoring discrepancies (low ROUGE-L but high BERTScore) for 4-HGRWL summaries.
@@ F.2 DECODA Discrepancy Data Points
Table 16 displays English translations of DECODA cases showing the most significant scoring discrepancies (low ROUGE-L but high BERTScore) for 4-HGRWL summaries.
@@
Tables 17 and 18 present original French DECODA cases with the most significant scoring discrepancies (low ROUGE-L but high BERTScore) for 4-WL and 4-HGRWL summaries, respectively, as analyzed in Section 4.3.
@@ Appendix G Datasets: License or Terms of Use
The DialogSum corpus used in this study comprises resources that are freely available online without copyright restrictions for academic use. For the DECODA dataset, the training and validation sets can be downloaded from their website131313https://pageperso.lis-lab.fr/benoit.favre/cccs/ upon acceptance of the corresponding usage and sharing terms, while the test set is available only upon request from the authors.